40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 4 12:46:27 2018, gautam, Update, ALS, Fiber ALS assay FiberALS.pdfFiberALS.svg.zip
    Reply  Thu Jan 11 14:22:40 2018, gautam, Update, ALS, Fiber ALS assay 
       Reply  Thu Jan 11 20:51:20 2018, gautam, Update, ALS, Fiber ALS assay IMG_5112.JPG
          Reply  Tue Jan 16 21:50:53 2018, gautam, Update, ALS, Fiber ALS assay 
             Reply  Thu Jan 18 00:35:00 2018, gautam, Update, ALS, Fiber ALS assay IR_ALS_20180118.pdfC2B4C1DD-6528-4067-9C13-6BD248629AD6.MOV
                Reply  Fri Jan 19 11:34:21 2018, gautam, Update, ALS, Fiber ALS assay IR_ALS_20180119.pdf
                   Reply  Fri Jan 19 23:04:11 2018, gautam, Update, ALS, Fiber ALS assay 
                      Reply  Wed Jan 24 00:33:31 2018, gautam, Update, ALS, Fiber ALS assay IMG_6866.JPG
                         Reply  Wed Jan 24 10:45:14 2018, gautam, Update, ALS, Fiber ALS assay 
                         Reply  Thu Jan 25 23:59:14 2018, gautam, Update, ALS, Fiber ALS assay 
                            Reply  Fri Jan 26 20:03:09 2018, gautam, Update, ALS, Fiber ALS assay 
                               Reply  Wed Jan 31 15:45:22 2018, gautam, Update, ALS, Fiber ALS assay IMG_6875.JPG
                                  Reply  Wed Jan 31 16:29:42 2018, gautam, Update, ALS, Modulation depths PMCmodDepth.pdfXPDH.pdf
                                  Reply  Wed Jan 31 16:33:53 2018, gautam, Update, ALS, ALS electronics at LSC rack 
                                     Reply  Wed Jan 31 22:32:11 2018, gautam, Update, ALS, ALS signal chain + power budget FiberALS_PowerBudget.pdfpreampProposed.pdf
                                        Reply  Thu Feb 1 01:24:56 2018, gautam, Update, ALS, D0902745 revamp underway PSLinterlock.png
                                           Reply  Fri Feb 2 00:26:34 2018, gautam, Update, ALS, D0902745 revamp underway 
                                              Reply  Fri Feb 2 13:16:55 2018, gautam, Update, ALS, ALS signals whitening switching ALS_whitening_switching.pdfALS_whitening_switching_works.pdf
                                                 Reply  Mon Feb 5 14:11:01 2018, gautam, Update, ALS, Huge harmonics in ALS channels ALSpowerlineHarmonics.pdf2A1B3AC4-4059-416A-AD88-8AB0431D7A55.jpeg
                                                    Reply  Mon Feb 5 22:57:28 2018, gautam, Update, ALS, Huge harmonics in ALS channels 
                                                       Reply  Tue Feb 6 11:13:26 2018, gautam, Update, ALS, Possible source of ground loop identified A68AF89C-E8A9-416D-BBD2-A1AD0A51E0B5.jpeg
                                                          Reply  Tue Feb 6 22:55:51 2018, gautam, Update, ALS, Possible source of ground loop identified 
                                                       Reply  Thu Feb 8 00:33:20 2018, gautam, Update, ALS, D990694 characterization / THD measurement plan D990694THD_trial.pdf
                                                          Reply  Thu Feb 8 01:27:16 2018, Koji, Update, ALS, D990694 characterization / THD measurement plan 
                                                          Reply  Fri Feb 9 13:37:44 2018, gautam, Update, ALS, THD measurement trial THD_trial.pdf
                                                       Reply  Thu Feb 8 12:00:09 2018, gautam, Update, ALS, D990694 is NOT differential receiving 
                                                          Reply  Thu Feb 8 13:13:14 2018, gautam, Update, ALS, D990694 pulled out IMG_5131.JPGIMG_5133.JPGD990694-B_cuttingPlan.pdf
                                                             Reply  Thu Feb 8 18:10:36 2018, gautam, Update, ALS, D990694 pulled out 
                                              Reply  Wed Feb 7 15:51:15 2018, gautam, Update, ALS, D0902745 revamp complete 6613CD37-5014-44AE-B1FE-6F6A8137DF62.jpeg
                                                 Reply  Fri Feb 16 01:34:40 2018, gautam, Update, ALS, D0902745 in-situ testing BeatMouthX_20180216.pdfIMG_5134.JPG
                                                    Reply  Tue Feb 20 23:08:27 2018, gautam, Update, ALS, D0902745 in-situ testing BeatMouthX_20180220_diffOut.pdf
                                                       Reply  Thu Feb 22 00:09:11 2018, gautam, Update, ALS, D0902745 in-situ testing daughterBoard_TF.pdfdaughterBoard_noise.pdfIMG_6906.JPG
                                                          Reply  Sun Feb 25 00:03:12 2018, gautam, Update, ALS, Daughter board prototyping schematic.pdfdaughterBoard_TF.pdfdaughterBoard_noise.pdfTEK00000.PNGdaughterBoard_noise.pdf
                                                             Reply  Mon Feb 26 20:55:56 2018, rana, Update, ALS, Daughter board prototyping 
                                                                Reply  Tue Feb 27 21:10:45 2018, gautam, Update, ALS, Daughter board testing IMG_6912.JPGIMG_6913.JPGIMG_6914.JPG
                                                                   Reply  Wed Feb 28 19:13:25 2018, gautam, Update, ALS, ADC test for differential receiving in c1lsc c1lsc_ADCtest.png
                                                                   Reply  Fri Mar 2 01:45:06 2018, gautam, Update, ALS, new look ALS electronics IMG_6916.JPG
                                                                      Reply  Mon Mar 5 17:27:34 2018, gautam, Update, ALS, new look ALS electronics - characterization BeatMouthX_20180305_diffOut.pdf
                                                                         Reply  Thu Mar 8 00:40:25 2018, gautam, Update, ALS, new look ALS electronics - characterization BeatMouthX_20180305_diffOut.pdf
                                                                            Reply  Thu Mar 8 19:38:37 2018, gautam, Update, ALS, digital unwhitening of daughter board TFfit.pdf
                                                                               Reply  Thu Mar 8 23:50:27 2018, gautam, Update, ALS, First look at new ALS electronics BeatMouth_OOL.pdf
                                                                                  Reply  Fri Mar 9 01:07:01 2018, gautam, Update, ALS, First look at new ALS electronics ALS_vs_POXnoise.pdf
                                                                                     Reply  Mon Mar 12 22:08:31 2018, gautam, Update, ALS, Noisy POX NoisyPOX_20180312.pdf
                                                                                        Reply  Mon Mar 12 23:57:31 2018, gautam, Update, ALS, Noisy POX NoisyPOXandPOY_20180312.pdfnoisyPOX_Jan2018.pdf
                                                                                           Reply  Mon Mar 19 15:02:29 2018, gautam, Update, ALS, Noisy MC sensing 
                                        Reply  Thu Feb 1 15:31:12 2018, gautam, Update, ALS, ALS signal chain + power budget 
                                        Reply  Wed Feb 7 10:16:26 2018, gautam, Update, ALS, ALS signal chain + power budget 
                      Reply  Thu Jan 25 13:18:41 2018, gautam, Update, ALS, Fiber ALS assay preampProposed.pdf
Message ID: 13636     Entry time: Fri Feb 16 01:34:40 2018     In reply to: 13616     Reply to this: 13644
Author: gautam 
Type: Update 
Category: ALS 
Subject: D0902745 in-situ testing 

Having implemented the changes to the audio amplifier stage, I re-installed this unit at the LSC rack, and did some testing. The motivation was to determine the shape of the ALS error signal spectrum, so that I can design a whitening preamp accordingly. Attachment #1 is the measurement I've been after. The measurement was taken with EX NPRO PDH locked to the arm via green, and Xarm locked to MC via POX. Slow temperature relief servo for EX NPRO was ON. Here are the details:

  1. Mode-matching into the BeatMouth PSL light fiber had deteriorated dramatically - it was ~1mW out of 4.4mW. I spent 5 mins getting it back to 3.2mW (72% efficiency) and then moved on... I am a little surprised the drift was so large, but perhaps, it's not surprising given that there has been a lot of work on and around the PSL table in the last couple of weeks. There is a 300mm focusing lens after the last steering mirror so the effect of any alignment drifts should be attenuated, I don't really understand why this happened. Anyways, perhaps a more intelligent telescope design would avoid this sort of problem.
  2. I removed the ND filter in the PSL pickoff to BeatMouth path (this was not responsible for the reduced power mentioned in #1). I verified that the total power reaching the photodiode was well below its rated damage threshold of 2mW (right now, there is ~620uW). I will update the BeatMouth schematic accordingly, but I think there will be more changes as we improve mode matching into the fibers at the end.
  3. Hooked up the output of the fiber PD to the Teledyne amp, routed the latters output to the LSC rack. Measured RF electrical power at various places. In summary, ~6dBm of beat reaches the splitter at the LSC rack. This is plenty.
  4. The main finding tonight was discovered by accident.
    • For the longest time, I was scratching my head over why the beat note amplitude, as monitored on the control room SA (I restored it to the control room from under the ITMX optical table where Koji had temporarily stored it for his tests on the PSL table) was drifting by ~10-15dB!
    • So each time, having convinced myself that the power levels made sense, I would come back to the control room to make a measurement, but then would see the beat signal level fluctuate slowly but with considerable amplitudeindecision.
    • The cause - See Attachment #2. There is a length of fiber on the PSL table that is unshielded to the BeatMouth. While plugging in RF cables to the BeatMouth, I found that accidentally brushing the fiber lightly with my arm dramatically changed the beat amplitude as monitored on a scope.
    • For now, I've "strain relieved" this fiber as best as I could, we should really fix this in a better way. This observation leads me to suspect that many of the peaky features seen in Attachment #1 are actually coupling in at this same fiber...
    • The beat note amplitude has been stable since, in the ~90 mins while I've been making plots/elogs.
    • Surely this is a consequence of differential polarization drift between the PSL and EX beams?
  5. There are prominent powerline harmonics in these signals - how can we eliminate these? The transmission line from PSL table to LSC rack already has a BALUN at its output to connect the signal to the unbalanced input of the demod board.
  6. Not sure what to make of the numerous peaks in the LO driven, RF terminated trace.
  7. The location of the lowest point in the bucket also doesn't quite match previous measured out-of-loop ALS noise - we seem to have the lowest frequency noise at 150-200Hz, but in these plots its more like 400Hz.

Conclusion: In the current configuration, with x10 gain on the demodulated signals, we barely have SNR of 10 at ~500Hz. I think the generic whitening scheme of 2 zeros @15Hz, 2poles@150Hz will work just fine. The point is to integrate this whitening with the preamp stage, so we can just go straight into an AA board and then the ADC (sending this signal into D990694 and doing the whitening there won't help with the SNR). Next task is to construct a test daughter board that can do this...

 

Attachment 1: BeatMouthX_20180216.pdf  289 kB  | Hide | Hide all
BeatMouthX_20180216.pdf
Attachment 2: IMG_5134.JPG  7.018 MB  | Hide | Hide all
IMG_5134.JPG
ELOG V3.1.3-